THE AGES OF THREE BROTHERS ARE CONSECUTIVE *EVEN* INTEGERS. THE SUM OF THEIR AGES IS 54. HOW OLD IS EACH OF THE BROTHERS?
Found 2 solutions by stanbon, haileytucki:
Answer by stanbon(75887)

THE AGES OF THREE BROTHERS ARE CONSECUTIVE *EVEN* INTEGERS. THE SUM OF THEIR AGES IS 54. HOW OLD IS EACH OF THE BROTHERS?
----------------------
1st: 2x-2
2nd: 2x
3rd: 2x+2
----
Equation:
sum = 54
6x = 54
---
x = 9
-----
1st: 2x-2 = 16
2nd: 2x = 18
3rd: 2x+2 = 20
==================
Cheers,
Stan H.

Answer by haileytucki(390)

54=x+(x+2)+(x+4)
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
x+(x+2)+(x+4)=54
Remove the parentheses that are not needed from the expression.
x+x+2+x+4=54
Since x and x are like terms, add x to x to get 2x.
2x+2+x+4=54
Since 2x and x are like terms, add x to 2x to get 3x.
3x+2+4=54
Add 4 to 2 to get 6.
3x+6=54
Since 6 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 6 from both sides.
3x=-6+54
Add 54 to -6 to get 48.
3x=48
Divide each term in the equation by 3.
(3x)/(3)=(48)/(3)
Simplify the left-hand side of the equation by canceling the common factors.
x=(48)/(3)
Simplify the right-hand side of the equation by simplifying each term.
x=16
16, 18, and 20